If Sarah is 24 years younger than her mother and if the sum of their ages is 68, how old is Sarah? whats an equation i can use for this
Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)

m is Mother's Age
m + (m-24) = 68
2m = 92
m = 46, Mother's Age. Sarah is 22 46-24
And...
46 + 22 = 68

Answer by Alan3354(69443)

M - S = 24
M + S = 68
--------------- Add